OpenStudy (anonymous):

Write a point-slope equation for the line with the given slope that contains the given point... m=-3(-2,-5)

OpenStudy (anonymous):

y=-3x-11

OpenStudy (chaise):

The straight line equation is: y=mx+b You're given your x coordinate, y coordinate and also your gradient. You are able to solve for b in this equation. -5 = -3x-2 + b -5 = 6 + b b = -11 y = -3x-11 Does this make sense? Do you want me to explain it further?

OpenStudy (anonymous):

Yes because my professor said this put what you know into the formula, y-y1=m(x-x1). You know a point which is an x1 & y1 and the slope, m - so put them in

OpenStudy (chaise):

One again; my enter button doesn't work. Sorry if this is hard to understand. Does anyone know how to fix this?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

Exactly what?

OpenStudy (anonymous):

wat u said

OpenStudy (anonymous):

I know but he wants me to put them into the formula

OpenStudy (anonymous):

well then u do becoz u have ur coordinates as well as ur m

OpenStudy (anonymous):

okay thanks